OpenStudy (anonymous):

Joan ate 100 grapes in five days, each day eating six more thank on the previous day. How many grapes did she eat on the first day? Please SHOW YOUR WORK

OpenStudy (anonymous):

6+12+18+24+30 + 500 lol then divide by 5 right?

hero (hero):

I was just going to ask where did bainsvsme get the plus 500? Either way, the appraoach is all wrong

OpenStudy (gw2011):

Let x=Number of grapes eaten on the first day 2nd day=6 more 3rd day=12 more 4th day=18 more 5th day=24 more Since x is involved each day, you get 5x 5x+6+12+18+24=100 5x+60=100 5x=100-60 5x=40 x=8 This is the number of grapes eaten on the first day